OpenStudy (anonymous):

i need help with this question, Find the 11th term in this sequence 1, -5, 25, -125,... and i found that the 11th term is -9,765,625 and now I need to know what the explicit formula is

OpenStudy (mathmath333):

this is a geometric sequence as the common ratio is constant \(\large \begin{align} \color{black}{ \normalsize \text{formula for nth term is } \hspace{.33em}\\~\\ n^{th} ~~term=a\times r^{n-1} \hspace{.33em}\\~\\ \normalsize \text{where a = first term , r = common ratio, n =nth term } \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

thanks man

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ 11^{th} ~~term=1\times (-5)^{11-1}=9765652 \hspace{.33em}\\~\\ }\end{align}\) not that u have put negative symbol before 9765652 which is wrong

OpenStudy (mathmath333):

\[\Huge \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#E6E600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#E6E600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{o}}\color{purple}{\text{p}}\color{purple}{\text{e}}\color{red}{\text{n}}\color{orange}{\text{s}}\color{#E6E600}{\text{t}}\color{green}{\text{u}}\color{blue}{\text{d}}\color{purple}{\text{y}}\color{purple}{\text{ }}\color{red}{\text{!}}\color{orange}{\text{}}\end{array} \]